A speeder traveling at 110 km/h races past a police car. The police car is at rest and pursues at acceleration of 1.25 m/s^2. How long does it take the cop car to catch the speeder answer in units of seconds? How fast is the police car going when it catches up to the speeder? Is this a realistic speed?
1
Expert's answer
2019-03-05T09:40:08-0500

We have


"vt=(at^2)\/2"

"v=at\/2"


So,


"t=2v\/a=2(110\/3.6)\/1.25=48.9 s."

The speed of the police car:


"V=at=a*2v\/a=2v=2(110\/3.6)=61.1 m\/( s)=220 km\/h."

It is a realistic speed.